If the pediatric dosage of chlorothiazide (DIURIL) is 10 to 20 mg/kg of body weight per day in a single dose or two divided doses, not to exceed 375 mg/day, calculate the daily dosage range of an oral suspension containing 250 mg chlorothiazide per 5 mL that should be administered to a 48-lb child.
Answer #1

Dosage of chlorthiazide – 10 to 20mg/kg/day in single or divided doses not exceeding 375mg/kg/day

1 lb = 2.2kg

48 lb = 48/2.2 = 21.8kg

Dosage for 48 lb child

10x21.8=218mg

20x21.8 = 436mg

19x21.8 = 414mg

18x21.8 = 392mg

17x21.8 = 370.6mg

Therefore the dosage = 218 to 370.6mg/day in single or divided doses

Dosage:

Dose ordered/ Dose on hand X Amount on hand = Amount to administer

218/250 X5 = 4.3

370.6/ 250 X5 = 7.4

Therefore the dosage of oral suspension = 4.3ml to 7.4ml / day in single or divided doses
Answer #2
A starting pediatric dose of phenytoin so- dium (DILANTIN Sodium) is 6 mg/kg/ d, administered in three equally divided doses. Using tablets containing 50 mg of phenytoin sodium, a pharmacist pre- pared a suspension such that each 1 mL, delivered from a calibrated dropper, con- tained a single dose for a 44-lb. child. How many tablets should be used to pre- pare 30 mL of the suspension?
Answer #3

Dosage of chlorthiazide – 10 to 20mg/kg/day in single or divided doses not exceeding 375mg/kg/day

1 lb = 2.2kg

48 lb = 48/2.2 = 21.8kg

Dosage for 48 lb child

10x21.8=218mg

20x21.8 = 436mg

19x21.8 = 414mg

18x21.8 = 392mg

17x21.8 = 370.6mg

Therefore the dosage = 218 to 370.6mg/day in single or divided doses

Dosage:

Dose ordered/ Dose on hand X Amount on hand = Amount to administer

218/250 X5 = 4.3

370.6/ 250 X5 = 7.4

Therefore the dosage of oral suspension = 4.3ml to 7.4ml / day in single or divided dose


answered by: JEEEEEEEEEEE